OpenStudy (anonymous):

Please solve: x+y=8. xy=15.

OpenStudy (anonymous):

x=8-y y(8-y)=15 0=y²-8y+15 0=(y-3)(y-5) that mean y = 3 or 5 when y = 3, x=8-(3)=5 when y=5 x=8-5=3